# YEAST: Yet Another Sequential Test

### v4

@@ 4.1 Type-I Error (False Detection Rate)
Two observations from our experiments are in order. Firstly, the proposed method (YEAST) with the robust variance estimator was the only method to demonstrate accurate false detection rate control (i.e. with the detection rate staying within one percentage point of the nominal significance level). Secondly, when the non-robust estimator was used, the detection rate was considerably inflated for all but one method-dataset combinations. This highlights the importance of accounting for the correlation in the events generated by the same user when estimating the variance.
@@ Appendix A Proofs and Auxiliary Statements
$\Pr(A_{\tau})=\Pr(A_{\bar{\tau}})\quad\forall\,\tau\in\mathcal{T}.$ (9)
Finally, fix an arbitrary Borel set $Q\subset\mathrm{I}\!\mathrm{R}$ and set
@@
Denoting the conditional probability $\Pr(\cdot\mid S_{1},\,\ldots,\,S_{k})$ by $\Pr_{k}$, we will now derive the key relationship for the proof of the lemma. Specifically, (3) and (9) imply that for any $\tau\in\mathcal{T}$ we have
@@
for all sufficiently large $N$. (12), relations (14)—-(15) give
@@ B.2 Non-Normal Data
The results are depicted in Figure 1. The first data point on each line corresponds to the case of no treatment effect and therefore represets the false detection rate. The remaining points represent the power for different treatment effect sizes. Similarly to the experiment with normal data, YEAST had a considerably higher power curve (both for the Gamma and the Student’s t distribution cases).
@@ Appendix C Discrete Monitoring
We report the share of replications where the test detects an effect (i. e. the significance is detected in at least one of the interim checks). Table 5 (“False Positive Rate”) reports this share for the case where no actual effect was present. All of the compared methods except CAA and oversampled versions of GST keep the false detection rate below the nominal level of 5%. Table 6 (“Power”) contains the share of replications with a detection for the case where the treatment effect was set to 0.2 standard deviations. The methods with an inflated false detection rate were excluded from the power comparisons. The GST method with cubic alpha-spending demonstrated the highest power, closely followed by YEAST and GST with quadratic alpha-spending. We find it remarkable that our proposed method, despite supporting continuous monitoring, performed on par with the GST method in the discrete monitoring setting.
@@ Appendix D Sample/Time Savings
Table 7: Simulation Experiment:
Sample/Time Savings
@@
1 YEAST 0.13 0.39 0.58 0.69
2 mSPRTphi100 0.09 0.35 0.62 0.75
3 mSPRTphi11 0.12 0.41 0.68 0.82
4 mSPRTphi25 0.12 0.41 0.68 0.81
5 GAVI250 0.12 0.40 0.67 0.80
6 GAVI500 0.10 0.37 0.63 0.77
7 GAVI750 0.08 0.34 0.60 0.74
8 Bonferroni 0.02 0.19 0.45 0.67
9 LanDeMetsOBF 0.14 0.41 0.60 0.70
10 SeqC2ST-QDA 0.03 0.06 0.14 0.27

### v5

@@ 4.1 Type-I Error (False Detection Rate)
Two observations from our experiments are in order. Firstly, the proposed method (YEAST) with the robust variance estimator was the only method to demonstrate accurate false detection rate control (i.e. with the detection rate staying within one percentage point of the nominal significance level). Secondly, when the non-robust estimator was used, the detection rate was considerably inflated. This highlights the importance of accounting for the correlation in the events generated by the same user when estimating the variance.
@@ Appendix A Proofs and Auxiliary Statements
$\Pr(A_{\tau}\mid Y_{1},\,W_{1},\,\ldots,\,Y_{k},\,W_{k})=\Pr(A_{\bar{\tau}}\mid Y_{1},\,W_{1},\,\ldots,\,Y_{k},\,W_{k})\quad\textup{a.s.}\quad\forall\,\tau\in\mathcal{T}.$ (9)
Finally, fix an arbitrary Borel set $Q\subset\mathrm{I}\!\mathrm{R}$ and define
@@
Denoting the conditional probability $\Pr(\cdot\mid Y_{1},\,W_{1},\,\ldots,\,Y_{k},\,W_{k})$ by $\Pr_{k}$, we will now derive the key relationship for the proof of the lemma. Specifically, (3) and (9) imply that for any $\tau\in\mathcal{T}$ we have [^5] In the second and next to the last transitions, we leverage the fact that if two collections of random variables are independent, they remain conditionally independent after conditioning on any subset of variables from either or both collections.
@@
which by the tower property implies that
$\Pr(\bar{B}\mid S_{1},\,\ldots,\,S_{k})=\Pr(B\mid S_{1},\,\ldots,\,S_{k})\quad\textup{a.s.}$ (11)
@@
for all sufficiently large $N$. (12), relations (14)–(15) give
@@ B.2 Non-Normal Data
The results are depicted in Figure 1. The first data point on each line corresponds to the case of no treatment effect and therefore represets the false detection rate. The remaining points represent the power for different treatment effect sizes. Similarly to the experiment with normal data, YEAST had a considerably higher power curve (both for the Gamma and the Student’s t distribution cases) than other methods that did not inflate the false detection rate.
@@ Appendix C Discrete Monitoring
We report the share of replications where the test detects an effect (i. e. the significance is detected in at least one of the interim checks). Table 5 (“False Positive Rate”) reports this share for the case where no actual effect was present. All of the compared methods except oversampled versions of GST keep the false detection rate below the nominal level of 5%. Table 6 (“Power”) contains the share of replications with a detection for the case where the treatment effect was set to 0.2 standard deviations. The methods with an inflated false detection rate were excluded from the power comparisons. The GST method with cubic alpha-spending demonstrated the highest power, closely followed by YEAST and GST with quadratic alpha-spending. We find it remarkable that our proposed method, despite supporting continuous monitoring, performed on par with the GST method in the discrete monitoring setting.
@@ Appendix D Sample/Time Savings
Table 7: Simulation Experiment: Sample/Time Savings, %
@@
1 YEAST 13 39 58 69
2 mSPRTphi100 9 35 62 75
3 mSPRTphi11 12 41 68 82
4 mSPRTphi25 12 41 68 81
5 GAVI250 12 40 67 80
6 GAVI500 10 37 63 77
7 GAVI750 8 34 60 74
8 Bonferroni 2 19 45 67
9 LanDeMetsOBF 14 41 60 70
10 SeqC2ST-QDA 3 6 14 27